A small park is in the shape of a rectangle that measures 525 feet by 468 feet. Fine the number of feet that would be needed to enclose the park. Express your answer to the nearest foot.
Answer by checkley77(12844)

2*525+2*468=1,050+936=1,986 feet will be need ed to enclose the park.